The brick wall (of thermal conductivity 1 W/m · ◦ C) of a building has dimensions of 4.5 m by 4 m and is 19 cm thick. How much heat flows through the wall in a 6.4 h period when the average inside and outside temperatures are, respectively, 16◦C and 2◦C? Answer in units of MJ.


1
Expert's answer
2021-05-26T10:33:46-0400

"Q=\lambda\frac{\Delta T}{\Delta x}At=1\cdot\frac{16-2}{0.19}\cdot(4.5\cdot4)\cdot(6.4\cdot3600)=30.56\ (MJ)" . Answer